An average human weighs about 600 N If two such g
Answer #1

F=K*q1q2/r^2

600 = 9*10^9*2*-2/r^2

600 = -36*10^9/r^2

r^2 = -36*10^9/600

r = sqrt(60*10^-6)

r = 7 745.966